# Pruning vs Quantization: Which is Better?

### v1

@@ Abstract
Neural network pruning and quantization techniques are almost as old as neural networks themselves. However, to date only ad-hoc comparisons between the two have been published. In this paper, we set out to answer the question on which is better: neural network quantization or pruning? By answering this question, we hope to inform design decisions made on neural network hardware going forward. We provide an extensive comparison between the two techniques for compressing deep neural networks. First, we give an analytical comparison of expected quantization and pruning error for general data distributions. Then, we provide lower bounds for the per-layer pruning and quantization error in trained networks, and compare these to empirical error after optimization. Finally, we provide an extensive experimental comparison for training 8 large-scale models on 3 tasks. Our results show that in most cases quantization outperforms pruning. Only in some scenarios with very high compression ratio, pruning might be beneficial from an accuracy standpoint.
@@ 3.4 Experiments on real weight tensors
The previous discussion was mostly theoretical. We set out to see happens when we do a similar analysis on real neural network weights. In order to investigate this, we compare the pruning and quantization SNR on the weight tensors for all the pre-trained models from the PyTorch model zoo111 https://pytorch.org/serve/model_zoo.html. (46 models in total, the details are give in appendix E). Each tensor is quantized using an integer grid of bit widths from 2 to 8. The results are shown in the figure 3 (left). We see a similar trend to our previous discussion that pruning becomes more beneficial for lower bitwidth/higher sparsity ratios.
@@ 5 Full-model comparison
In our experiments we used a set of 8 models trained for 3 tasks including Resnet18, Resnet50 [25], MobileNet-V2 [54], MobileNet-V3-small [27], EfficientNet-lite [56], and ViT [10] trained on ImageNet classification [53]; DeepLab-V3 [6] with MobileNet-V2 backbone trained for semantic segmentation on Pascal VOC [12]; EfficientDet [57] trained for object detection on MS COCO [40].
@@
| pruning | 34.5 | 33.0 | 30.9 | 27.9 | 24.2 | 17.9 | 8.0 |
@@ 6 Discussion
To provide insights into representations learned during pruning or QAT, we studied the evolution of models during fine-tuning. We found that fine-tuning after pruning tends to recover the original representation, while quantization-aware training leads to learning completely new representations. We provide further details on these experiments in appendix H.
@@
Impact Using pruning or quantization leads to power reduction on many architectures and enables new applications on mobile platforms. We see only a positive impact from this on the whole.
@@ 7 Related work
Despite the large amount of work on pruning, quantization, and combining them, there is little literature comparing the two methods. To the best of our knowledge, there is only one work that performs a comparison of pruning versus non-uniform quantization [30]. The work considers only small-scale models and provides only an empirical comparison with no further analysis.
@@ Appendix G Details of the full-model experiments
Appendix H Analysis of representations learned during fine-tuning in QAT and pruning

### v2

@@ Abstract
Neural network pruning and quantization techniques are almost as old as neural networks themselves. However, to date only ad-hoc comparisons between the two have been published. In this paper, we set out to answer the question on which is better: neural network quantization or pruning? By answering this question, we hope to inform design decisions made on neural network hardware going forward. We provide an extensive comparison between the two techniques for compressing deep neural networks. First, we give an analytical comparison of expected quantization and pruning error for general data distributions. Then, we provide lower bounds for the per-layer pruning and quantization error in trained networks, and compare these to empirical error after optimization. Finally, we provide an extensive experimental comparison for training 9 large-scale models on 4 tasks. Our results show that in most cases quantization outperforms pruning. Only in some scenarios with very high compression ratio, pruning might be beneficial from an accuracy standpoint. 111Code is available at https://github.com/Qualcomm-AI-research/pruning-vs-quantization
@@ 3.4 Experiments on real weight tensors
The previous discussion was mostly theoretical. We set out to see happens when we do a similar analysis on real neural network weights. In order to investigate this, we compare the pruning and quantization SNR on the weight tensors for all the pre-trained models from the PyTorch model zoo111 https://pytorch.org/serve/model_zoo.html. (46 models in total, the details are give in appendix E) combined with weight tensors from 3 large language models, namely Bloom-3b [3], Llama-3b [20], OPT-2.7b [67]. Each tensor is quantized using an integer grid of bit widths from 2 to 8. The results are shown in the figure 3(left). We see a similar trend to our previous discussion that pruning becomes more beneficial for lower bitwidth/higher sparsity ratios.
@@ 5 Full-model comparison
In our experiments we used a set of 4 models trained for 4 tasks including Resnet18, Resnet50 [27], MobileNet-V2 [58], MobileNet-V3-small [30], EfficientNet-lite [60], and ViT [11] trained on ImageNet classification [57]; DeepLab-V3 [7] with MobileNet-V2 backbone trained for semantic segmentation on Pascal VOC [13]; EfficientDet [61] trained for object detection on MS COCO [43]; OPT-350 fine-tuned on WikiText-103.
@@
To study the effect of training time, we also performed an ablation with 2 times longer fine-tuning on a subset of 3 models (Resnet50, EfficientNet, and ViT). The results are given in appendix H. We observe that pruned models generally benefit from fine-tuning more, and in particular pruning becomes more beneficial for most compression ratios on Resnet50. However, for the other models, quantization is still more beneficial due to a larger gap in performance.
@@
pruning | 34.5 | 33.0 | 30.9 | 27.9 | 24.2 | 17.9 | 8.0 | |||
OPT-350m | 14.8 | perpl. | quant. | 14.8 | 14.8 | 14.9 | 15.0 | 15.3 | 15.9 | 19.9 |
pruning | 18.0 | 19.7 | 22.6 | 27.2 | 35.4 | 53.5 | 101.4 |
Combining pruning and quantization
Another interesting questions is whether pruning is beneficial in combination with quantization. To answer it, we perfromed an experiment on pruning quantized Resnet-18, MobileNet-V2 and ViT with different pruning ratios. The results are given on figure 5. On x-axis we plot the expected bit-widths which is a product of the base bit-width and the sparsity in the pruned model including the natural sparsity. The points marked by crosses are quantized models with only natural sparsity and no extra pruning applied. As we can see, mild degrees of pruning are beneficial in the combinations. However, we note that no extra overhead was assumed for storing the pruning mask.
@@ 6 Discussion
To provide insights into representations learned during pruning or QAT, we studied the evolution of models during fine-tuning. We found that fine-tuning after pruning tends to recover the original representation, while quantization-aware training leads to learning completely new representations. We provide further details on these experiments in appendix I.
@@
Impact Using pruning or quantization leads to power reduction on many architectures and enables new applications on mobile platforms. We see only a positive impact from this on the whole. In some cases both pruning and quantization might lead to biased predictions, a further discussion can be found in [29].
@@ 7 Related work
Despite the large amount of work on pruning, quantization, and combining them, there is little literature comparing the two methods. To the best of our knowledge, the closest work that performs a comparison of pruning versus non-uniform quantization [33]. The work considers only small-scale models and provides only an empirical comparison with no further analysis. Another related study is [53].
@@ 8 Conclusion
9 Acknowledgement
We would like to thank Marios Fournarakis and Yelisei Bondarenko for their help with performing QAT experiments.
@@ Appendix F Details of per-layer experiments
To reduce the computational complexity of finding the global solution for pruning, the layers had to be split into chunks. The slice of 4 input channels over all output channels was used for 3x3 convolutions. In the case of linear layers and point-wise convolutions, slices 36 input features over all the output features were used.
@@ Appendix G Details of the full-model experiments
In our experiments we used a set of 9 models trained for 4 tasks including Resnet18, Resnet50 [27], MobileNet-V2 [58], MobileNet-V3-small [30], EfficientNet-lite [60], and ViT [11] trained on ImageNet classification [57]; DeepLab-V3 [7] with MobileNet-V2 backbone trained for semantic segmentation on Pascal VOC [13]; EfficientDet [61] trained for object detection on MS COCO [43], and OPT-350m fine-tuned on WikiText-103.
@@
Appendix H Full-model experiments with longer fine-tuning
In this section we report the results for QAT and magnitude pruning with twice as longer fine-tuning compared to table 4. The results are given in table 5
Model | Orig. | Metric | Method | 8b | 7b | 6b | 5b | 4b | 3b | 2b |
Resnet-50 | 76.1 | acc. | quant. | 76.6 | 76.6 | 76.5 | 76.3 | 76.2 | 75.5 | 72.3 |
pruning | 76.6 | 76.7 | 76.6 | 76.3 | 76.3 | 75.8 | 74.8 | |||
EfficientNet | 75.4 | acc. | quant. | 75.2 | 75.3 | 75.0 | 74.6 | 74.0 | 71.8 | 61.5 |
pruning | 73.0 | 71.9 | 69.5 | 65.8 | 60.0 | 50.7 | 34.6 | |||
ViT | 81.3 | acc. | quant. | 81.9 | 81.8 | 81.7 | 81.4 | 80.8 | 78.9 | 73.7 |
pruning | 80.3 | 79.7 | 79.0 | 77.7 | 76.3 | 74.3 | 71.6 |
Appendix I Analysis of representations learned during fine-tuning in QAT and pruning